OpenStudy (anonymous):

Find V1 using mesh and nodal analysis

OpenStudy (anonymous):

Here is the Figure

OpenStudy (mostadorthsander):

are there any links as to how to interpret the components? what is 8i1...is that a fpga or a plc or something like that??

OpenStudy (anonymous):

its the current flowing on 2 ohms resistor

OpenStudy (anonymous):

If I interpret your diagram correct: let top left node be Va, top right Vb, mid left Vd, mid right Vc, ground reference V=0 at bottom. Vb-Va = 8i1 Va+16 = Vd Vd = -i1*2 (-Va/4) + (-Vd/2) + (-Vc/2) = 3V1/4 Vc - Vb = V1 5 = V1/2 + Vc/4 Looks like 6 equations, 6 unknowns.